# VisualMimic Visual Humanoid Loco-Manipulation via Motion Tracking and Generation

### v1

@@ I Introduction
Equipping humanoid robots with such human-like object interaction abilities has been a long-standing challenge. Current approaches can be categorized into three main paradigms based on tasks: First, locomotion-focused methods [1, 2] that excel at terrain traversal but do not address object interaction. Second, approaches that rely on external motion capture systems [3, 4] for object state estimation, restricting their deployment to controlled laboratory environments. Third, vision-based methods for object interaction, which follow two distinct paths: 1) imitation learning approaches [5, 6, 7] that train visuomotor policies via human demonstrations, which are constrained by the scarcity of large-scale demonstration data and result in limited generalization capabilities; and 2) sim-to-real reinforcement learning (RL) methods [8, 9] that offer greater robustness and generalizability; however, vision-based RL is currently limited to simple environmental interactions such as sitting [8] and stair climbing [8, 9], falling significantly short of human-level object interaction abilities, due to the large exploration and action space of humanoid robots.
@@ II Related Work
TWIST [5]
VideoMimic [8]
Hitter [3]
Recipe [10]
HEAD [11]
@@ II-A Learning Humanoid Loco-Manipulation
Enabling humanoid robots to perform versatile loco-manipulation in unstructured environments, akin to humans, is a long-standing goal for roboticists. Currently, two main pathways are being explored: (1) Imitation learning on real-world data collected via whole-body teleoperation [6, 7, 5, 12, 13]. While these methods demonstrate promising task versatility, they remain limited by the scarcity of high-quality data and the difficulty of scaling data collection. (2) Sim-to-real reinforcement learning based on large-scale simulation interaction [8, 3, 14, 15, 9, 16]. These approaches exhibit strong generalization in specific humanoid motor skills, such as terrain traversal [9, 15], box picking [16], and table tennis [3], but are restricted in task diversity compared to imitation learning. Several works remain confined to simulation; for example, HumanoidBench [14] and SkillBlender [17] adopt hierarchical frameworks similar to ours. However, their policies are often excessively jittery or depend on privileged object states, hindering successful deployment in the real world.
Previous works on learning visuomotor policies for humanoid robots have typically focused either on upper-body manipulation [18, 19, 20, 10] or on perceptive locomotion [9, 21, 22]. More recently, VideoMimic [8] introduced a real2sim2real pipeline that enables real robots to perform environment interactions such as sitting, though their interactions remain limited to static settings like the ground or stone chairs. Other efforts, such as PDC [23], have demonstrated promising results in vision-based whole-body loco-manipulation, but only in simulation. In contrast, we propose a sim-to-real framework that enables real-world humanoid robots to achieve versatile object interaction and loco-manipulation through egocentric vision.
@@ III-A General Keypoint Tracker
Although a keypoint tracking policy can be trained directly, its ability to capture motion is weaker than that of a motion tracking policy due to the much simplified commands, resulting in less human-like behaviors (Fig. 6). We address this problem by designing a two-stage teacher-student distillation pipeline. Specifically, in the teacher training stage, a privileged teacher motion tracker is trained via RL with access to future reference motions. Then, a student keypoint tracker is trained using DAgger [24], relying solely on proprioception and keypoint commands computed from the reference frame at each time step.
@@
Since the teacher policy is not used during deployment, we provide it with sufficient motion and proprioceptive information so that it can track as accurately as possible. The teacher motion tracker $\pi_{\text{tracker}}^{\text{tea}}$ takes as input a sequence of future reference motion frames (spanning 2 seconds) and privileged proprioceptive information (e.g., foot contact forces), which allows it to anticipate upcoming goals and generate smoother motion. We implement $\pi_{\text{tracker}}^{\text{tea}}$ as a simple three-layer MLP and optimize it using PPO [25, 26]. Following the reward structure in [5], our reward $r_{\text{motion}}$ encourages accurate motion tracking while penalizing artifacts such as jitter and foot slippage: $r_{\text{motion}}=r_{\text{track}}+r_{\text{penalty}}$. Notably, we track robot body positions and root velocities in the world frame. For motion datasets, we use GMR [27, 5] to retarget AMASS [28] and OMOMO [29] into humanoid motions.
@@
After obtaining $\pi_{\text{generator}}^{\text{tea}}$, we distill it into $\pi_{\text{generator}}^{\text{stu}}$ via DAgger [24], which takes keypoint commands $c^{\text{kp}}_{t}$ as input and is real-world deployable. We define
@@ III-B Task-Specific Keypoint Generator
Directly training such a keypoint generator with visual RL, however, is highly inefficient (Table III), because crucial information—such as object positions and contact forces—is only partially observable, and incorporating vision into IsaacGym further slows down the simulation. To overcome these challenges, we adopt a two-stage approach to train a task-specific keypoint generator. In the first stage, a teacher generator with access to task-relevant object states is trained using PPO [25]. In the second stage, a student keypoint tracker that relies only on depth images for object information is distilled.
@@
As discussed in the previous paragraph, the teacher keypoint generator leverages object states to accelerate training. The object state is defined within the environment and concatenated with proprioceptive information to form the input of the state-based keypoint generator. The generator, implemented as a three-layer MLP, is trained using PPO [25, 26] with a task-specific reward function.
@@
Since object states are unavailable during deployment, the state-based keypoint generator $\pi_{\text{generator}}^{\text{tea}}$ is distilled into a student keypoint generator that relies solely on visual observations and proprioceptive inputs. Since RGB images have significant sim-to-real gaps, we use depth images as the only visual modality. The depth input is processed by a CNN encoder, whose output is then concatenated with proprioceptive features and passed through an MLP. The student keypoint generator is distilled using DAgger [24].

### v2

@@ I Introduction
Equipping humanoid robots with such human-like object interaction abilities has been a long-standing challenge. Current approaches can be categorized into three main paradigms based on tasks: First, locomotion-focused methods [RealHumanoid2023, TokenHumanoid2024] that excel at terrain traversal but do not address object interaction. Second, approaches that rely on external motion capture systems [su2025hitter, xue2025full] for object state estimation, restricting their deployment to controlled laboratory environments. Third, vision-based methods for object interaction, which follow two distinct paths: 1) imitation learning approaches [ze2025twist, ben2024homie, li2025amo] that train visuomotor policies via human demonstrations, which are constrained by the scarcity of large-scale demonstration data and result in limited generalization capabilities; and 2) sim-to-real reinforcement learning (RL) methods [videomimic, wang2025beamdojo] that offer greater robustness and generalizability; however, vision-based RL is currently limited to simple environmental interactions such as sitting [videomimic] and stair climbing [videomimic, wang2025beamdojo], falling significantly short of human-level object interaction abilities, due to the large exploration and action space of humanoid robots.
@@ II Related Work
TWIST [ze2025twist]
VideoMimic [videomimic]
Hitter [su2025hitter]
Recipe [lin2025sim]
HEAD [head]
@@ II-A Learning Humanoid Loco-Manipulation
Enabling humanoid robots to perform versatile loco-manipulation in unstructured environments, akin to humans, is a long-standing goal for roboticists. Currently, two main pathways are being explored: (1) Imitation learning on real-world data collected via whole-body teleoperation [ben2024homie, li2025amo, ze2025twist, he2024omnih2o, fu2024humanplus]. While these methods demonstrate promising task versatility, they remain limited by the scarcity of high-quality data and the difficulty of scaling data collection. (2) Sim-to-real reinforcement learning based on large-scale simulation interaction [videomimic, su2025hitter, sferrazza2024humanoidbench, chen2024lcp, wang2025beamdojo, dao2024sim]. These approaches exhibit strong generalization in specific humanoid motor skills, such as terrain traversal [wang2025beamdojo, chen2024lcp], box picking [dao2024sim], and table tennis [su2025hitter], but are restricted in task diversity compared to imitation learning. Several works remain confined to simulation; for example, HumanoidBench [sferrazza2024humanoidbench] and SkillBlender [kuang2025skillblender] adopt hierarchical frameworks similar to ours. However, their policies are often excessively jittery or depend on privileged object states, hindering successful deployment in the real world.
Previous works on learning visuomotor policies for humanoid robots have typically focused either on upper-body manipulation [ze2024idp3, cheng2024tv, qiu2025humanoid, lin2025sim] or on perceptive locomotion [wang2025beamdojo, He2025AttentionBased, long2024learning]. More recently, VideoMimic [videomimic] introduced a real2sim2real pipeline that enables real robots to perform environment interactions such as sitting, though their interactions remain limited to static settings like the ground or stone chairs. Other efforts, such as PDC [luo2025emergent], have demonstrated promising results in vision-based whole-body loco-manipulation, but only in simulation. In contrast, we propose a sim-to-real framework that enables real-world humanoid robots to achieve versatile object interaction and loco-manipulation through egocentric vision.
@@ III-A General Keypoint Tracker
Although a keypoint tracking policy can be trained directly, its ability to capture motion is weaker than that of a motion tracking policy due to the much simplified commands, resulting in less human-like behaviors (Fig. 6). We address this problem by designing a two-stage teacher-student distillation pipeline. Specifically, in the teacher training stage, a privileged teacher motion tracker is trained via RL with access to future reference motions. Then, a student keypoint tracker is trained using DAgger [ross2011reduction], relying solely on proprioception and keypoint commands computed from the reference frame at each time step.
@@
Since the teacher policy is not used during deployment, we provide it with sufficient motion and proprioceptive information so that it can track as accurately as possible. The teacher motion tracker $\pi_{\text{tracker}}^{\text{tea}}$ takes as input a sequence of future reference motion frames (spanning 2 seconds) and privileged proprioceptive information (e.g., foot contact forces), which allows it to anticipate upcoming goals and generate smoother motion. We implement $\pi_{\text{tracker}}^{\text{tea}}$ as a simple three-layer MLP and optimize it using PPO [schulman2017ppo, rudin2022legged_gym]. Following the reward structure in [ze2025twist], our reward $r_{\text{motion}}$ encourages accurate motion tracking while penalizing artifacts such as jitter and foot slippage: $r_{\text{motion}}=r_{\text{track}}+r_{\text{penalty}}$. Notably, we track robot body positions and root velocities in the world frame. For motion datasets, we use GMR [ze2025gmr, ze2025twist] to retarget AMASS [mahmood2019amass] and OMOMO [li2023omomo] into humanoid motions.
@@
After obtaining $\pi_{\text{generator}}^{\text{tea}}$, we distill it into $\pi_{\text{generator}}^{\text{stu}}$ via DAgger [ross2011reduction], which takes keypoint commands $c^{\text{kp}}_{t}$ as input and is real-world deployable. We define
@@ III-B Task-Specific Keypoint Generator
Directly training such a keypoint generator with visual RL, however, is highly inefficient (Table III), because crucial information—such as object positions and contact forces—is only partially observable, and incorporating vision into IsaacGym further slows down the simulation. To overcome these challenges, we adopt a two-stage approach to train a task-specific keypoint generator. In the first stage, a teacher generator with access to task-relevant object states is trained using PPO [schulman2017ppo]. In the second stage, a student keypoint tracker that relies only on depth images for object information is distilled.
@@
As discussed in the previous paragraph, the teacher keypoint generator leverages object states to accelerate training. The object state is defined within the environment and concatenated with proprioceptive information to form the input of the state-based keypoint generator. The generator, implemented as a three-layer MLP, is trained using PPO [schulman2017ppo, rudin2022legged_gym] with a task-specific reward function.
@@
Since object states are unavailable during deployment, the state-based keypoint generator $\pi_{\text{generator}}^{\text{tea}}$ is distilled into a student keypoint generator that relies solely on visual observations and proprioceptive inputs. Since RGB images have significant sim-to-real gaps, we use depth images as the only visual modality. The depth input is processed by a CNN encoder, whose output is then concatenated with proprioceptive features and passed through an MLP. The student keypoint generator is distilled using DAgger [ross2011reduction].